# Symmetry-aware spectral bounds for the electronic structure Hamiltonian

### v1

@@ Abstract
We present symmetry-aware spectral bounds to assess the query complexity of Hamiltonian oracles in quantum algorithms. Our numerical estimates indicate that these bounds are smaller than traditional ones for a variety of electronic structure systems, including exhibiting unique scaling behavior in thermodynamic and complete basis set limits. Our work highlights potential room for improvement in reducing the $\ell_{1}$ norm through tensor factorization and block-encoding methods, while also offering a critical evaluation of query complexity limits for quantum algorithms in physics and chemistry.
@@
Quantum computing promises to play a pivotal role in physics and chemistry, targeting key problem domains like quantum simulation, Gibbs state preparation, solving linear systems and partial differential equations, as well as precise energy and observable estimation [1, 2, 3, 4, 5, 6, 7, 8, 9, 4, 10]. In recent years, a variety of quantum algorithms have emerged to tackle these problems, each offering distinct trade-offs in runtime and resource requirements [1, 2, 3, 11, 12, 13, 14, 15, 16]. At the heart of these approaches are Hamiltonian oracles, such as Hamiltonian evolution and block-encoding oracles, which are critical in assessing a quantum algorithm’s total resource requirements.
@@
The computational cost of quantum algorithms for physics and chemistry is often determined by the spectrum of the Hamiltonian. Many algorithms require a rescaling of the Hamiltonian to ensure that its output eigenvalues are confined within certain operational bounds, impacting the precision needed for each algorithm and thus influencing the query complexity of Hamiltonian oracles. In quantum phase estimation and its modern variants, this scaling is crucial for mitigating aliasing and potential spectral leakage effects [17, 18]. The scaling also emerges naturally within the linear combination of unitaries (LCU) or block-encoding methods to ensure the oracle remains unitary [19]. The need for normalization or rescaling transcends specific algorithmic implementations and is a universal requirement across various quantum computing paradigms. Appendix A provides a brief overview and summary of how Hamiltonian rescaling affects the query complexity of quantum algorithms.
@@
As a concrete example within the LCU and block-encoding framework, QPE based on qubitization has become a leading fault-tolerant algorithm for estimating the ground-state energy [2, 20, 21, 22]. To achieve a desired precision $\epsilon$, excluding the initial state overlap dependence, the total number of Toffoli gates is known to follow the asymptotic scaling, $\mathcal{O}\!\left(\lceil\tfrac{\lambda}{\epsilon}\rceil C_{\mathcal{W}[H]}\right)$, where $\lambda$ is equal to the $\ell_{1}$ norm required for block-encoding the Hamiltonian, while $C_{\mathcal{W}[H]}$ denotes the total number of Toffoli gates needed to compile the qubitization walk operator $\mathcal{W}[H]$. In this context, $\lceil\frac{\lambda}{\epsilon}\rceil$ specifies the number of query calls, which is rounded up to the nearest integer or power of two [22]. This scaling dependence highlights that while reducing $\lambda$ lowers query complexity, reducing $C_{\mathcal{W}[H]}$ decreases the oracle gate complexity.
@@
Within the context of electronic structure theory, these concepts have led to the development of tensor factorization and compression techniques, which aim to provide efficient representations of the Hamiltonian [23, 21, 22, 24, 25]. These methods have a long history in classical algorithms for quantum chemistry, where approaches like density fitting and Cholesky decompositions have long been used [26, 27, 28, 29], while tensor hypercontraction (THC) techniques have been developed over the past fifteen years to provide further enhancements [30, 31]. In quantum computing, methods such as Double Factorization, Regularized Compressed Double Factorization (RCDF), as well as THC have shown great promise in reducing the resource requirements of quantum algorithms [21, 23, 32, 22, 33, 24, 25]. These techniques have proven effective in reducing the computational cost by either minimizing the $\ell_{1}$ norm $\lambda$, which impacts query complexity, or by decreasing the total number of terms in the Hamiltonian, affecting $C_{\mathcal{W}[H]}$.
As the development of tensor compression methods continues, it is essential to understand their fundamental limits. Accurate lower bounds for $\lambda$ and $C_{\mathcal{W}[H]}$ would establish the minimal resource cost for many quantum algorithms based on block-encoding, irrespective of the specific tensor compression scheme. To this end, the recent work by Loaiza et al. [34] derived the spectral bound:
@@
For a bounded Hermitian operator $\hat{H}$, all of its possible LCU decompositions have an associated $\ell_{1}$ norm $\lambda$ which is lower bounded by the spectral range, $\Delta/2\equiv(E_{\rm max}-E_{\rm min})/2$,
@@
While this theorem is significant for bounded Hermitian operators, it does not take into account prior knowledge of potential symmetries. This suggests the spectral bound might overestimate the spectral range for symmetry sectors relevant to many problem instances.
@@
In this work, we present a corollary and additional theorems that leverage the symmetries of the Hermitian operator $\hat{H}$ to establish symmetry-aware lower bounds on the $\ell_{1}$ norm. We also propose a variational and numerically tractable approach for computing lower bounds of symmetry-sector spectral ranges, showing that numerical estimates of symmetry-sector bounds can be much smaller than previous bounds for many problem instances relevant to electronic structure theory. Our work analyzes the unique scaling behavior of these bounds in the thermodynamic and complete basis set limits for various molecular systems, while also highlighting potential room for improvement in tensor factorization and block-encoding methods. The aim of this work is to provide a better understanding of the query complexity limits for different Hamiltonian oracles in physics and chemistry.
@@
Symmetry-aware spectral bounds. To take into account the effect of symmetries, let us define the symmetry transformation generated by an operator, $\hat{U}_{S}$, that leaves the Hamiltonian unchanged such that $\hat{U}_{S}^{\dagger}\hat{H}\hat{U}_{S}=\hat{H}\rightarrow[\hat{H},\hat{U}_{S}]=0$. It is then possible to show that the Hamiltonian $\hat{H}$ can be rotated into a block-diagonal basis, $\hat{H}=\bigoplus_{\mu}\hat{H}_{\mu},$ where $\hat{H}_{\mu}$ is a sub-block defined with respect to a particular symmetry sector $\mu$. This observation leads to the following important corollary, which describes two additional lower bounds:
@@
The proofs of Theorem 1 and Corollary 2 are provided in Appendix B. While $\Delta/2$ provides a lower bound on the $\ell_{1}$ norm $\lambda$ of the original Hamiltonian, Theorem 3 in Appendix B shows that $\Delta_{s}/2$ is the lower bound of the $\ell_{1}$ norm of symmetry-shifted Hamiltonian proposed in [34]. Furthermore, Theorem 4 in Appendix B also proves that $\Delta_{\mu}/2$ lower bounds the so-called block-invariant Hamiltonian proposed in [35]. This substantiates the prevailing intuition that the symmetry-sector spectral range should inherently determine the query complexity lower bound for Hamiltonian oracles. Given that all three Hamiltonians share an equivalent eigenvalue spectrum within the symmetry sector $\mu$, they are equally viable for constructing Hamiltonian oracles tailored to problem instances confined to this sector. These results not only set lower bounds on the query complexity for such oracles but also provide a hard cut-off in the expected improvement of tensor compression strategies designed to minimize the $\ell_{1}$ norm.

### v2

# Assessing the query complexity limits of quantum phase estimation using symmetry aware spectral bounds

@@ Abstract
The computational cost of quantum algorithms for physics and chemistry is closely linked to the spectrum of the Hamiltonian, a property that manifests in the necessary rescaling of its eigenvalues. The typical approach of using the 1-norm as an upper bound to the spectral norm to rescale the Hamiltonian suits the most general case of bounded Hermitian operators but neglects the influence of symmetries commonly found in chemical systems. In this work, we introduce a hierarchy of symmetry-aware spectral bounds that provide a unified understanding of the performance of quantum phase estimation algorithms using block-encoded electronic structure Hamiltonians. We present a variational and numerically tractable method for computing these bounds, based on orbital optimization, to demonstrate that the computed bounds are smaller than conventional spectral bounds for a variety of molecular benchmark systems. We also highlight the unique analytical and numerical scaling behavior of these bounds in the thermodynamic and complete basis set limits. Our work shows that there is room for improvement in reducing the 1-norm, not yet achieved through methods like double factorization and tensor hypercontraction, but highlights potential challenges in improving the performance of current quantum algorithms beyond small constant factors through 1-norm reduction techniques alone.
I Introduction
@@
Quantum computing holds the potential to significantly impact physics and chemistry, addressing various problem areas such as quantum simulation, Gibbs state sampling, solving linear systems and partial differential equations, as well as precise energy and observable estimation [1, 2, 3, 4, 5, 6, 7, 8, 9, 4, 10]. Over the past few years, a wide range of quantum algorithms have emerged, each offering distinct trade-offs in runtime and resource cost [1, 2, 3, 11, 12, 13, 14, 15, 16].
@@
The computational cost of quantum algorithms is intricately linked to the spectrum of the Hamiltonian. Many algorithms require a rescaling of the Hamiltonian to ensure that its output eigenvalues are confined within certain operational bounds, impacting the precision needed for each algorithm and thus influencing the query complexity of Hamiltonian oracles. In quantum phase estimation and its modern variants [13, 14, 15, 16], this is crucial for mitigating aliasing and potential spectral leakage effects [17, 18]. The scaling also emerges naturally within linear combination of unitaries (LCU) or block-encoding methods to ensure the oracle remains unitary [19]. The need for rescaling often transcends specific algorithms and is required across various quantum computing paradigms.
@@
For example, within the context of block-encoding and the LCU framework, qubitization-based QPE stands out as a leading fault-tolerant algorithm for estimating the ground-state energy [2, 20, 21, 22]. To achieve a desired precision $\epsilon$, excluding the initial state overlap dependence, the total number of Toffoli gates is known to follow the asymptotic scaling, $\mathcal{O}\!\left(\lceil\tfrac{\lambda}{\epsilon}\rceil C_{\mathcal{W}[H]}\right)$, where $\lambda$ is equal to the 1-norm required for block-encoding the Hamiltonian, $C_{\mathcal{W}[H]}$ denotes the total number of Toffoli gates needed to compile the qubitization walk operator $\mathcal{W}[H]$, and $\lceil\frac{\lambda}{\epsilon}\rceil$ specifies the number of query calls, which is rounded up to the nearest integer or power of two [22]. This scaling dependence highlights that reducing both $\lambda$ and $C_{\mathcal{W}[H]}$ is necessary for improving the total gate complexity of this quantum algorithm.
@@
In electronic structure theory, these concepts have led to the development of tensor factorization techniques, which aim to provide efficient representations of the Hamiltonian [23, 21, 22, 24, 25]. For context, these methods have a long history in the development of classical algorithms for quantum chemistry, where approaches like density fitting and Cholesky decompositions are now well established [26, 27, 28, 29]. Techniques like tensor hypercontraction (THC) have also been developed over the past fifteen years to provide further enhancements [30, 31]. In quantum computing, methods like double factorization, regularized compressed double factorization, and THC have been used to reduce the resource cost of quantum algorithms [21, 23, 32, 22, 33, 24, 25], by minimizing either the 1-norm $\lambda$, or qubitization oracle complexity, $C_{\mathcal{W}[H]}$. It is unclear, however, whether these techniques are close to optimal, or whether there remains ample room for improvement.
Our work contributes to understanding the fundamental limits of these tensor factorization techniques both theoretically and numerically. Specifically, we study the query complexity limits arising from lower bounds on the 1-norm $\lambda$, deferring the examination of $C_{\mathcal{W}[H]}$ for future research. To this end, we build upon the findings of Loaiza et al. [34], who established a spectral bound for the 1-norm $\lambda$:
@@
For a bounded Hermitian operator $\hat{H}$, all of its possible LCU decompositions have an associated 1-norm $\lambda$ which is lower bounded by half the spectral range, $\Delta/2\equiv(E_{\rm max}-E_{\rm min})/2$,
@@
This Theorem highlights that the spectral range, $\Delta$, equal to the difference between the largest and smallest eigenvalues, acts as a robust lower bound to the 1-norm $\lambda$. The spectral range is invariant to the choice of unitary operators in the LCU decomposition, which distinguishes it from the spectral norm, $\|\hat{H}\|$, equal to the largest singular value of $\hat{H}$, which is sometimes used as a lower bound to the 1-norm. In particular scenarios, like the complete-square structure observed in double factorization [21, 34], the spectral norm can numerically overestimate the lowest achievable lower bound due to the presence of undesired identity terms that may present in the LCU decomposition.
Despite the significance of Theorem 1 for bounded Hermitian operators, two fundamental concerns emerge. First, the Theorem does not take into account prior knowledge of potential symmetries in the Hermitian operator, $\hat{H}$. This omission suggests that the spectral bound, $\Delta/2$, might overestimate the spectral range for symmetry sectors relevant to many Hamiltonians and problem instances. Second, the theorem does not offer a practical method for computing such bounds numerically. Past approaches relied on exact diagonalization methods, effective only for small-scale electronic structure systems expanded with minimal basis sets [34, 35]. Consequently, evaluating these bounds for large-scale systems like FeMoco or P450, which have become standard benchmark systems for quantum computing algorithms, remains challenging.
@@
In this work, we present a Corollary and additional Theorems that leverage the symmetries of the Hermitian operator $\hat{H}$ to establish symmetry-aware lower bounds to the 1-norm. We also propose a variational and numerically tractable approach for computing lower bounds of symmetry-sector spectral ranges, showing that numerical estimates of symmetry-sector spectral ranges can be significantly smaller than conventional bounds for different problem instances relevant to electronic structure theory. In addition, our work analyzes the unique scaling behavior of these bounds in the thermodynamic and complete basis set limits for various molecular systems. The aim of this work is to provide a better understanding of the query complexity limits for different Hamiltonian oracles commonly used in physics and chemistry, highlighting potential room for improvement, as well as future challenges that may arise in tensor factorization and block-encoding methods.
The paper is structured as follows. In Section II, we summarize several preliminary results and definitions that will be used throughout the manuscript. In section III, we present the main theoretical results of the paper concerning the hierarchy of symmetry-aware spectral bounds, as well as theorems that establish a connection between these bounds and corresponding 1-norms of various symmetry-shifted Hamiltonians. In section IV, we present the computational procedure used for computing these bounds numerically. In section V, the numerical results are presented for large-scale benchmark systems, such as FeMoco and P450. Finally, we present numerical evidence for the scaling behavior of the symmetry-aware spectral bounds in the thermodynamic and complete basis set limits.
II Preliminary Definitions
To encode the physical problem of interest, we define a Hamiltonian, $\hat{H}$, expressed as an LCU decomposition consisting of a sum of unitary operators $\hat{U}_{k}$ weighted by complex coefficients $c_{k}$,
$\hat{H}=\sum_{k}^{L}c_{k}\hat{U}_{k},\;\;\text{where}\;\;\lambda=\sum_{k}^{L}|c_{k}|,$ (2)
is the 1-norm of these coefficients. Throughout this document, we also impose the requirement that the Hamiltonian $\hat{H}$ is invariant under the action of a set of symmetries, $\mathbb{S}=\{\hat{S}_{m}\}$, which collectively form a group $G$, satisfying the conditions $[\hat{S}_{m},\hat{S}_{n}]=[\hat{S}_{m},\hat{H}]=0,\forall\;m,n\in 1,\cdots,|\mathbb{S}|$. The group $G$ includes all individual group elements for each symmetry, as well as any combinations of these symmetries that leave the Hamiltonian invariant. Maschke’s theorem enables a decomposition of the Hilbert space $\mathcal{H}$ into invariant subspaces, which are determined to be irreducible if they lack further invariant subspaces [43, 44]. Schur’s lemma may then be used to reveal the block-diagonal structure of $\hat{H}$ as,
$\hat{H}=\bigoplus_{\boldsymbol{m}}\bigoplus_{n=1}^{\mathrm{deg}\,\boldsymbol{m}}\hat{H}_{n}^{\boldsymbol{m}}\equiv\bigoplus_{\mu}\hat{H}_{\mu},$ (3)
In the following paragraphs, we discuss the importance of the Hamiltonian spectrum, and in particular the spectral range which acts as a key parameter, affecting the performance of a broad range of quantum algorithms that employ either the Hamiltonian evolution (HE) or block-encoding (BE) oracles. Both of these Hamiltonian oracles play a key role in a wide variety of quantum algorithms with applications in preparing zero-temperature ground states, preparing finite temperature Gibbs states, estimating Green’s functions and correlation functions, as well as energy and observable estimation.
Hamiltonian Evolution Oracle. The first oracle we consider is the Hamiltonian evolution model defined by,
$U_{\mathrm{HE}}=e^{i\hat{H}}.$ (4)
When using quantum phase estimation, this Hamiltonian oracle must be normalized so that the eigenvalue spectrum lies between 0 and $2\pi$ to avoid aliasing, specially relevant in the limit of arbitrary input states. This is achieved by using a modified Hamiltonian, $H^{\prime}=c_{1}H+c_{2}\mathbb{1}$, where $c_{1}$ and $c_{2}$ are normalization constants. The choice of these constants depends on the Hamiltonian type and the input state used in the algorithm. The constant $c_{1}$ is crucial as it rescales the eigenvalue spectrum, influencing the required precision for the output phase.
For general bounded Hermitian operators without symmetries, the ideal normalization constants are $c_{1}=\frac{2\pi}{\Delta}$ and $c_{2}=-\frac{2\pi}{\Delta}E_{\mathrm{min}}$, where $\Delta=E_{\mathrm{max}}-E_{\mathrm{min}}$ is the spectral range, and $E_{\mathrm{min}}/E_{\mathrm{max}}$ are the minimum/maximum eigenvalues of the Hamiltonian operator, $\hat{H}$. These normalization constants are suitable for arbitrary input states. Although these eigenvalues are not known in advance, it is possible to use approximate eigenvalues $\tilde{E}_{\mathrm{min}}/\tilde{E}_{\mathrm{max}}$ satisfying the conditions, $E_{\mathrm{min}}\leq\tilde{E}_{\mathrm{min}}$ and $E_{\mathrm{max}}\leq\tilde{E}_{\mathrm{max}}$. As an example, the methods presented in Appendix A and Section IV may be used to provide numerically tractable estimates of $\tilde{E}_{\mathrm{min}}/\tilde{E}_{\mathrm{max}}$.
On the other hand, if the Hamiltonian preserves a specific symmetry transformation, different rescaling parameters should be used. Assuming the input state to the QPE algorithm belongs to the symmetry sector $\mu$, the normalization constants should ideally become equal to $c_{1}=\frac{2\pi}{\Delta_{\mu}}$ and $c_{2}=-\frac{2\pi}{\Delta_{\mu}}E^{\mu}_{\mathrm{min}}$, where $\Delta_{\mu}=E^{\mu}_{\mathrm{max}}-E^{\mu}_{\mathrm{min}}$, and $E^{\mu}_{\mathrm{min}}/E^{\mu}_{\mathrm{max}}$ are the minimum/maximum eigenvalues of the Hamiltonian operator $\hat{H}$ in symmetry sector $\mu$ ($\hat{H}_{\mu}$). The rescaling constants are suitable for arbitrary input states belonging to symmetry sector $\mu$. As above, the methods presented in Appendix A and Section IV may be used to provide numerically tractable estimates of $\tilde{E}^{\mu}_{\mathrm{min}}/\tilde{E}^{\mu}_{\mathrm{max}}$ belonging to a given symmetry sector $\mu$.
Block-Encoding Oracle. The second oracle we consider is the block-encoding oracle model defined by,
$U_{\mathrm{BE}}=\begin{pmatrix}\hat{H}/\lambda&*\\ *&*\end{pmatrix}.$ (5)
Block-encoding oracles have similar applications to HE oracles since these two input models are in fact equivalent if we ignore the overhead that comes from converting between them. Within the context of ground-state energy estimation, block-encoding is often used as a building block of the qubitization walk operator, defined as $\mathcal{W}=\exp\left(\pm i\cos^{-1}(\hat{H}/\lambda)\right)$. Since the output phase, $\varphi=\pm\cos^{-1}(E/\lambda)$, is limited between 00 and $\pi$, aliasing is not a fundamental concern in this context. The primary consideration is ensuring that $\mathcal{W}$ remains unitary for any input state. This implies the Hamiltonian oracle must have an eigenvalue spectrum bounded between -1 and 1.
For general bounded Hermitian operators without any symmetries, the normalization constant $\lambda$ is equal to the 1-norm in Eqn (2). Since the 1-norm is highly dependent on the specifics of the Hamiltonian LCU expansion, including orbital optimization and tensor compression methods [38, 21, 22], it is possible to provide a rough lower bound using the spectral range, $\frac{\Delta}{2}\leq\|\hat{H}\|$, as highlighted in Theorem 1 [34]. This implies that the rescaling for general Hamiltonians in the block-encoding oracle model is also limited by the spectral range for arbitrary input states.
The situation changes when Hamiltonian preserves set of symmetries $\mathbb{S}=\{\hat{S}_{m}\}$. As highlighted in [34, 35], it is possible to define two new Hamiltonians, $H_{S}$, and $H_{BI}$, which themselves have different eigenvalue spectra from $\hat{H}$.
Definition II.1 (Symmetry-shifted Hamiltonian).
The symmetry-shifted Hamiltonian is defined as,
$\hat{H}_{\mathrm{S}}=\hat{H}-\hat{f}(\mathbb{S}),$ (6)
Definition II.2 (Block-invariant Hamiltonian).
The block-invariant Hamiltonian is defined as,
$\hat{H}_{\mathrm{BI}}=\hat{H}_{\mathrm{S}}-\hat{B}(\hat{f}(\mathbb{S})-f_{\mu}),$ (7)
where $\hat{B}$ is a Hermitian operator that obeys the same symmetries as the Hamiltonian, $[\hat{B},\hat{f}(\mathbb{S})]=0$, but it is assumed that it does not generally commute with the Hamiltonian. The operator $\hat{f}(\mathbb{S})$ is another well-behaved function of the Hamiltonian’s symmetries, $[\hat{H},\hat{f}(\mathbb{S})]=0$, obeying the eigenvalue relation, $\hat{f}(\mathbb{S})|\psi_{\mu}\rangle=f_{\mu}|\psi_{\mu}\rangle$.
While the symmetry-shifted Hamiltonian $\hat{H}_{S}$ preserves the eigenvalue spectrum for all possible symmetry-preserving input states apart from symmetry-sector-dependent identity shifts, the block-invariant Hamiltonian $\hat{H}_{\mathrm{BI}}$ preserves the correct eigenvalue spectrum for the symmetry sector $\mu$ only. As a result, it is possible to show that the symmetry sector bound acts as the true lower bound, $\frac{\Delta_{\mu}}{2}\leq\lambda_{\mathrm{BI}}$, for arbitrary input states. The details of this inequality are presented explicitly in the section below.
III Symmetry-aware spectral bounds
@@
To take into account the effect of symmetries, we use the results from the previous section and explicitly write the Hamiltonian $\hat{H}$ in the block-diagonal basis, $\hat{H}=\bigoplus_{\mu}\hat{H}_{\mu},$ where $\hat{H}_{\mu}$ is a sub-block defined with respect to a particular symmetry sector $\mu$. This leads to the following corollary, which describes the hierarchy of inequalities:
@@
Proof.
Note that the original spectral bound can be expressed as:
$\frac{\Delta}{2}=\frac{1}{2}\left(\max_{\mu}E_{\mathrm{max}}^{\mu}-\min_{\mu}E^{\mu}_{\mathrm{min}}\right).$ (9)
$\Delta$ $=\max_{\mu}E_{\mathrm{max}}^{\mu}-\min_{\mu}E^{\mu}_{\mathrm{min}}$ (10)
$=\max_{\mu}E_{\mathrm{max}}^{\mu}-\theta$ (11)
$=\max_{\mu}(E_{\mathrm{max}}^{\mu}-\theta)$ (12)
$\Delta_{s}\equiv\max_{\mu}(E_{\mathrm{max}}^{\mu}-E_{\mathrm{min}}^{\mu})\leq\Delta.$ (13)
While this Corollary proves that a hierarchy of spectral bounds exists, it does not prove that a symmetry-shifted $\hat{H}_{\mathrm{S}}$ or block-invariant $\hat{H}_{\mathrm{BI}}$ Hamiltonian must have corresponding 1-norms that are restricted by either of these two bounds. The following two theorems prove that these lower bounds apply directly to symmetry-shifted and block-invariant Hamiltonians introduced in refs. [34, 35].
Theorem 3 (Symmetry-shifted 1-norm bound).
For a bounded symmetry-shifted Hermitian operator, $\hat{H}_{\mathrm{S}}$, all of its possible LCU decompositions will have an associated 1-norm $\lambda_{\mathrm{S}}$ which is lower bounded by the spectral symmetry bound, $\Delta_{s}/2\equiv\tfrac{1}{2}\max_{\mu}(E_{\mathrm{max}}^{\mu}-E_{\mathrm{min}}^{\mu})$,
$\tfrac{1}{2}\Delta_{s}\leq\lambda_{\mathrm{S}}.$ (14)
Proof.
$\frac{1}{2}\left(\max_{\mu}(E_{\mathrm{max}}^{\mu}-f_{\mu})-\min_{\mu}(E^{\mu}_{\mathrm{min}}-f_{\mu})\right)\leq\lambda_{\mathrm{S}}.$ (15)
$\frac{1}{2}\left(\max_{\mu}(E_{\mathrm{max}}^{\mu}-f_{\mu})-\min_{\mu}(E^{\mu}_{\mathrm{min}}-f_{\mu})\right)$ $=\frac{1}{2}\left(\max_{\mu}(E_{\mathrm{max}}^{\mu}-f_{\mu})+\max_{\mu}(f_{\mu}-E^{\mu}_{\mathrm{min}})\right)$ (16)
$\geq\frac{1}{2}\max_{\mu}\left(E_{\mathrm{max}}^{\mu}-f_{\mu}+f_{\mu}-E^{\mu}_{\mathrm{min}}\right).$ (17)
Theorem 4 (Block-invariant 1-norm bound).
Consider the block-invariant Hermitian operator, $\hat{H}_{\mathrm{BI}}$, defined in (7) with an optimal symmetry-shifted Hamiltonian $\hat{H}_{\mathrm{S}}$. All of its possible LCU decompositions will have an associated 1-norm $\lambda_{\mathrm{BI}}$ which is lower bounded by the symmetry sector bound, $\Delta_{\mu}/2\equiv\tfrac{1}{2}(E_{\mathrm{max}}^{\mu}-E_{\mathrm{min}}^{\mu})$,
$\tfrac{1}{2}\Delta_{\mu}\leq\lambda_{\mathrm{BI}}.$ (18)
Proof.
The first term in (7) corresponds to the symmetry-shifted Hamiltonian. If we assume that the symmetry-shifted Hamiltonian has been optimized so that it saturates its bound, $\tfrac{1}{2}\Delta_{s}$, then it is possible to show that $\lambda_{\mathrm{BI}}$ can become smaller than $\tfrac{1}{2}\Delta_{s}$ for a certain set of operators $\hat{B}$ and $\hat{f}(S)$. This illustrates that $\tfrac{1}{2}\Delta_{s}$ is not a proper lower bound. Nevertheless, Corollary 2 shows that $\lambda_{\mathrm{BI}}$ must remain lower bounded by the symmetry sector bound, $\tfrac{1}{2}\Delta_{\mu}$, regardless of the chosen operators $\hat{B}$ and $\hat{f}(\mathbb{S})$. This shows that $\tfrac{1}{2}\Delta_{\mu}$ is the true lower bound of the 1-norm of the block-invariant Hamiltonian, thereby completing the proof. ∎
Ultimately, while the conventional spectral bound, $\Delta/2$, provides a lower bound on the 1-norm $\lambda$ of the original Hamiltonian, Theorem 3 shows that $\Delta_{s}/2$ is the lower bound of the 1-norm of symmetry-shifted Hamiltonian proposed in [34]. Furthermore, Theorem 4 proves that $\Delta_{\mu}/2$ lower bounds the so-called block-invariant Hamiltonian proposed in [35]. This substantiates the prevailing intuition that the symmetry-sector spectral range should inherently determine the query complexity lower bound for Hamiltonian oracles in quantum phase estimation. Given that all three Hamiltonians share an equivalent eigenvalue spectrum within the symmetry sector $\mu$, they are equally viable for constructing Hamiltonian oracles tailored to problem instances confined to this sector. These results not only set lower bounds on the query complexity for such oracles but also provide a hard cut-off in the expected improvement of tensor compression strategies designed to minimize the 1-norm.
III.1 Electronic Structure Hamiltonian